A man is able to save #50 of his salary in a particular year ,after this,every year he saved #20 more than the preceeding year .how long does it take him to save #4370
Answer by mananth(16946)


A man is able to save #50 of his salary in a particular year ,after this,every year he saved #20 more than the preceeding year .how long does it take him to save #4370
First year 50
seond year 70
third year 90
time = n = ?
Sn = 4370
Sn= [ (n/2)(2a+(n-1)d)]
4370 = n/2[ 2*50 + (n-1)*20]
8740= n(100+20n-20)
8740 = n(80+20n)
8740 = 80n +20n^2
20n^2+80n-8740=0
/20
n^2+4n -437=0
(n-19)(n+23) =0
Taking positive root
n=19
19 years